### v1

# Crowd-PrefRL: Preference-Based Reward Learning from Crowds

Abstract
Preference-based reinforcement learning (RL) provides a framework to train agents using human feedback through pairwise preferences over pairs of behaviors, enabling agents to learn desired behaviors when it is difficult to specify a numerical reward function. While this paradigm leverages human feedback, it currently treats the feedback as given by a single human user. Meanwhile, incorporating preference feedback from crowds (i.e. ensembles of users) in a robust manner remains a challenge, and the problem of training RL agents using feedback from multiple human users remains understudied. In this work, we introduce Crowd-PrefRL, a framework for performing preference-based RL leveraging feedback from crowds. This work demonstrates the viability of learning reward functions from preference feedback provided by crowds of unknown expertise and reliability. Crowd-PrefRL not only robustly aggregates the crowd preference feedback, but also estimates the reliability of each user within the crowd using only the (noisy) crowdsourced preference comparisons. Most importantly, we show that agents trained with Crowd-PrefRL outperform agents trained with majority-vote preferences or preferences from any individual user in most cases, especially when the spread of user error rates among the crowd is large. Results further suggest that our method can identify minority viewpoints within the crowd.
1 Introduction
Figure 1: Crowd-PrefRL framework for preference-based RL from crowds. We assume a crowd, or ensemble, of users is queried for their preferences over pairs of segmented behaviors (A vs. B). These crowd preferences are then used to learn a reward function for preference-based RL.
Reinforcement learning (RL) from human feedback Casper et al. (2023) is a promising approach for learning intelligent behaviors in the absence of a known numerical reward signal. Such methods have shown success in domains such as Atari games Christiano et al. (2017), robotics Wilde et al. (2021); Torne et al. (2023); Lee et al. (2021), and large language models (LLMs) Ouyang et al. (2022). In this work, we consider human feedback that is crowdsourced, i.e., where (1) data is labeled by multiple human users (e.g. $\geq 2$ labels for each data point), and (2) these labels are aggregated to form an ensemble label to be used for downstream analysis. In applications from LLMs to household robotics to autonomous driving, crowdsourced human data could enable RL algorithms to learn user-preferred behaviors that are robust across many real-world situations. Yet, at the same time, crowdsourcing algorithms will need to recognize and account for disagreement across users in the crowd. Example scenarios could include: an algorithm that a) trains models that are sensitive to multiple user viewpoints (e.g., an LLM provides information that benefits an underrepresented minority at minimal cost to the majority Siththaranjan, Laidlaw, and Hadfield-Menell (2023)), b) trains models that reflect a majority viewpoint while enabling downstream individual personalization (e.g., a household robot demonstrates performant out-of-the-box behavior, but fine-tunes to specific users’ preferences for tidying and organization), and c) identifies and rejects malicious user data (e.g., an LLM is resilient to feedback containing misinformation).
This work considers RL from crowdsourced pairwise preferences, in which a human answers queries of the form “Do you prefer A or B?” Christiano et al. (2017); Lee, Smith, and Abbeel (2021); Lee et al. (2021). Current preference-based RL methods treat preference feedback as given by a single human user, even if the feedback is actually given by multiple users with potentially different backgrounds and levels of expertise. For example, when learning reward functions for training LLMs, each individual preference pair query is typically shown to just one of the users due to multiple constraints (e.g. cost, human labeling time/effort); thus, the full labeled dataset aggregates the disjoint query subsets shown to each user Ouyang et al. (2022). However, it may often be preferable to not only obtain feedback from a group of users, but also to aggregate the crowdsourced data in a non-naive manner that recognizes differences between individuals. This capability can be beneficial for multiple reasons, for example to analyze patterns of disagreement in preference feedback across the crowd and to detect minority groups in the crowd (i.e., users that provide feedback that is consistently different from the majority). It also holds important practical modeling implications, for example mitigating model bias and ensuring fairness Ouyang et al. (2022) and assisting users with diverse preferences or varying degrees of noise in their feedback Siththaranjan, Laidlaw, and Hadfield-Menell (2023).
While previous literature uses crowdsourcing methods to aggregate labels across users and learn crowd preferences Chen et al. (2013); Li et al. (2016), the specific application of crowdsourcing to reward learning in RL remains understudied. Early works that explicitly model human feedback from crowds have focused on imitation learning, where there are multiple demonstrations (assumed to be from multiple human users) of unknown expertise and quality, and the goal is to learn a policy reflecting expert behavior Babes et al. (2011); Dimitrakakis and Rothkopf (2012); Hausman et al. (2017). As noted in Beliaev et al. (2022), however, these approaches simply aggregate the feedback from multiple humans into one homogeneous group, potentially leading to sub-optimal learning due to the presence of varying levels of expertise and reliability among the different human users. Recent work that has explicitly modeled crowdsourced data for imitation learning Beliaev et al. (2022) has shown significant performance gains from simultaneously modeling human demonstrations and estimating the demonstrators’ expertise levels. While recent work by  Zhang and Kashima (2023) used crowdsourcing methods to learn a reward function using preference-based reward learning, it is limited to the offline RL case Levine et al. (2020), where there is a fixed dataset of preference pairs, and the RL algorithm does not have online access to the environment. This potentially limits the expressiveness of the learned reward function to the behaviors contained in the offline dataset.
In this work, we study preference-based RL leveraging feedback from crowds in the online RL scenario, where the RL algorithm has access to the environment and can sample new preference pairs for the crowd to label throughout model training. Specifically:
  1. 1.
    We introduce the Crowd-PrefRL framework for performing online preference-based RL leveraging feedback from crowds and instantiate Crowd-PrefRL using Proximal Policy Optimization (PPO) Schulman et al. (2017), a state-of-the-art RL algorithm, to propose Crowd-PrefPPO. Crowd-PrefRL aggregates the user feedback to learn crowdsourced preference labels, and then uses those labels for reward function learning in preference-based RL (see Figure 1).
  2. 2.
    We use techniques derived from unsupervised ensemble learning to robustly aggregate the crowd preference feedback, estimating each user’s reliability in a completely unsupervised fashion and detecting the presence of multiple, e.g. minority, viewpoints among the crowd.
  3. 3.
    We perform experiments suggesting that RL agents trained with crowd preferences outperform RL agents trained via majority-vote preferences or preferences from any individual crowd member, in particular when the spread of user errors among the crowd is large. We also show a proof-of-concept for the ability of our method to identify minority viewpoints within a crowd.
This work demonstrates the viability of learning reward functions from preference feedback provided by crowds of unknown expertise and reliability. We believe that this is the first work that focuses on using preference feedback from crowds to learn a crowdsourced reward function for preference-based RL in the online RL setting.
2 Problem Formulation
In this paper, we are concerned with answering the following questions: (1) How can we effectively combine preference feedback from multiple users of unknown expertise and unknown reliability in the absence of ground-truth information? (2) Can we derive who are the “best” users in a given crowd in an unsupervised manner? (3) Can we learn a crowdsourced reward function that is better than any individual user’s reward function?, and (4) Can we detect the presence of minority feedback (as opposed to purely noisy feedback) in the crowd based only on the preference feedback? This section formulates this problem more precisely.
We consider an agent that takes actions and interacts with an environment Sutton and Barto (2018), such that at each timestep $t$, the agent receives a state $\mathbf{s}_{t}$ from the environment and chooses an action $\mathbf{a}_{t}$ based on its policy $\pi$. In traditional episodic RL, the environment also returns a reward $r(\mathbf{s}_{t},\mathbf{a}_{t})$, and the agent seeks to maximize the total discounted sum of rewards over an episode. However, for many complex tasks, it is difficult to construct a suitable reward function. Preference-based RL approaches instead query a human user to obtain preference feedback over segments of agent behaviors, and use this feedback to learn a reward function  Christiano et al. (2017); Lee, Smith, and Abbeel (2021). A segment is defined as a length-$H$ sequence of observations and actions, $\{(\mathbf{s}_{1},\mathbf{a}_{1}),...,(\mathbf{s}_{H},\mathbf{a}_{H})\}$. Given a pair of segments $x=(A,B)$, a user indicates which segment is preferred: $+1$ in the event that $A$ is preferred over $B$ (denoted by $A\succ B$) or $-1$ if $B$ is preferred over $A$ (denoted by $A\prec B$). Preference-based RL aims to train an agent to perform human-desirable behaviors according to the human user’s preferences.
Our problem formulation assumes access to a crowd, or ensemble, of users, who each assign pairwise preference labels to a sequentially-selected set $\mathcal{S}=\{x_{k}\}_{k=1}^{S}$, $S=|\mathcal{S}|$ of segment pairs. Each query $x_{k}$ takes the form $x_{k}=(A_{k},B_{k})$, where $A_{k},B_{k}$ belong to the set $\mathcal{X}$ of available segments. In addition, we assume that a majority (e.g. at least $51\%$) of the crowd provides preference feedback according to a shared understanding of desired agent behavior. More concretely, we assume the existence of an unknown ground-truth reward $r$, and that for a majority of users, their preferences are (noisily) generated with respect to $r$. Note that this formulation allows a minority group of users to potentially provide feedback according to a different reward function(s). Beyond this, we do not assume any further knowledge about the crowd (e.g. expertise, reliability). Our goal is to learn a crowd-informed preference label for each segment pair using only the preference labels given by the crowd. Formally, let ${\{f_{i}\}_{i=1}^{M}}$ represent $M$ users of unknown reliability and error rate, where each user $i$ provides a preference label $f_{i}(x_{k})\in\{-1,1\}$ on each input query $x_{k}\in\mathcal{S}$. Then, let $\bm{y}$ be the vector of (unknown to the algorithm) ground-truth labels according to the true reward $r$ (i.e., in which segments with higher ground-truth reward are preferred), $\bm{y}=[y_{1},...,y_{S}]^{T}$. Using only the preference labels from $M$ users on the segment pairs in $\mathcal{S}$ and without access to any ground-truth rewards or preferences, we seek an optimal estimate $\hat{\bm{y}}=[\hat{y}_{1},...,\hat{y}_{S}]^{T}$ of the true labels $\bm{y}$ given the crowd preference labels. We also investigate the ability to identify minority groups of users who provide feedback that consistently differs from that of the majority.
3 Methods
We next explain our methodology for simulating diverse crowds of human users with different rationality levels. Afterwards, we outline our Crowd-PrefRL approach to performing RL with crowdsourced feedback (Algorithm 1): first, we explain how to estimate preference labels from a crowd and then, we show how to integrate these crowd-aggregated preference labels into a framework for preference-based RL.
3.1 Simulating Diverse Crowds
In order to systematically evaluate and study the effects of user error rates and their impact on learning from crowdsourced user preferences, we apply the Stochastic Preference Model Lee et al. (2021) to simulate users with mixtures of different irrationality levels:
$P[A\succ B;\beta,\gamma]=$ (1)
$\hskip 14.22636pt\frac{e^{\left(\beta\sum_{t=1}^{H}\gamma^{H-t}r(\mathbf{s}_{t}^{i},\mathbf{a}_{t}^{i})\right)}}{e^{\left(\beta\sum_{t=1}^{H}\gamma^{H-t}r(\mathbf{s}_{t}^{i},\mathbf{a}_{t}^{i})\right)}+e^{\left(\beta\sum_{t=1}^{H}\gamma^{H-t}r(\mathbf{s}_{t}^{j},\mathbf{a}_{t}^{j})\right)}},$
where $\gamma\in(0,1]$ is a discount factor to model myopic behavior and $\beta$ is a rationality constant. Note that user feedback is fully rational and deterministic as $\beta\rightarrow\infty$, whereas $\beta=0$ produces uniformly random choices. In addition, we model an additional parameter $\epsilon$ that models users making mistakes due to accidental errors. For our analysis, each user is simulated with a random draw of parameters from the following ranges: $\gamma\in[0.96,1.0]$, $\beta\in[-0.5,-1.0]$, and $\epsilon\in[0.1,0.5]$. These parameter ranges where chosen as they produced diverse user error rates primarily between $15\%$ and $40\%$, which is approximately the range of human error rates observed in a prior study of human feedback in preference-based RL ($12\%-37\%$ in Brown et al. (2019)). We then sample crowds of $M\in\{7,11,15\}$ users according to this model to study the effect of crowd size on crowd label learning.
3.2 Estimating Preference Labels from a Crowd
For each query, Crowd-PrefRL distills the $M$ preference labels from each crowd member to a single preference label. We discuss two methods for accomplishing this distillation. The first, majority voting, is a comparatively naive approach, while the second, the Spectral Meta Learner (SML) Parisi et al. (2014) is the method we recommend for use with Crowd-PrefRL. Note that our Crowd-PrefRL framework generalizes the crowdsourcing approach in Brown et al. (2019), in which majority voting is used to merge preference labels.
Majority Vote (Baseline). Given the current problem formulation, where we do not know any additional information about the crowd, perhaps the simplest estimate $\hat{\bm{y}}$ is the majority vote (MAJ), which labels each segment pair $x_{k}\in\mathcal{S}$ with the majority preference from the crowd, denoted as $\hat{\bm{y}}_{MAJ}$.
Algorithm 1 Crowd-PrefRL: Preference-based RL with reward learning from crowds
1:frequency of crowd feedback $K$
2:number of queries $N_{\tt query}$ per feedback session
3:crowd size $M$
4:crowdsource method, one of (“MAJ”, “SML”)
5:Initialize parameters of policy $\pi_{\phi}$, reward model $\widehat{r}_{\psi}$, preference dataset $\mathcal{D}\leftarrow\emptyset$, and buffer $\mathcal{B}\leftarrow\emptyset$
6:// Exploration phase
7:$\mathcal{B},\pi_{\phi}\leftarrow\texttt{EXPLORE}()$ according to policy $\pi_{\phi}$
8:for each iteration do
9:     // Reward learning
10:     if iteration % $K==0$ then
11:         for $m$ in $1\ldots N_{\tt query}$ do
12:              Sample pair of segments $(A,B)$
13:              $\hat{\bm{y}}_{1:M}$leftarrow query all $M$ users in crowd
14:         end for
15:         if crowd method == “SML then
16:              Compute SML ranking vector $\hat{\bm{v}}_{i}$
17:              Compute $\hat{\bm{y}}_{SML}$ according to (2)
18:              $\hat{\bm{y}}\leftarrow\hat{\bm{y}}_{SML}$
19:         else Compute MAJ label estimate $\hat{\bm{y}}_{MAJ}$
20:              $\hat{\bm{y}}\leftarrow\hat{\bm{y}}_{MAJ}$
21:         end if
22:         $\mathcal{D}\leftarrow\mathcal{D}\cup\{(A,B,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\hat{\bm{y}}})\}$
23:         for each gradient step do
24:              Sample batch $\{(A,B,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\hat{\bm{y}}})_{j}\}_{j=1}^{|\mathcal{D}|}\sim\mathcal{D}$
25:              Optimize $\mathcal{L}^{\tt Reward}$ (4) with respect to $\psi$
26:         end for
27:     end if
28:     // Policy learning
29:     for each timestep $t$ do
30:         Collect $\mathbf{s}_{t+1}$ by taking $\mathbf{a}_{t}\sim\pi(\mathbf{a}_{t}|\mathbf{s}_{t})$
31:         $\mathcal{B}\leftarrow\mathcal{B}\cup\{(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1},\widehat{r}_{\psi}(\mathbf{s}_{t}))\}$
32:     end for
33:     for each gradient step do
34:         Sample random minibatch $\{\left(\tau_{j}\right)\}_{j=1}^{B}\sim\mathcal{B}$
35:         Optimize RL objective with respect to $\phi$
36:     end for
37:     Reset $\mathcal{B}\leftarrow\emptyset$ if on-policy RL algorithm is used
38:end for
Spectral Meta-Learner. However, it is indeed possible to estimate a crowd label that often outperforms the majority vote in the fully unsupervised case studied here. Under the assumption that users in a crowd make independent errors, Parisi et al. (2014) derived the Spectral Meta-Learner (SML):
$\hat{\bm{y}}_{k}=\operatorname{sign}\left(\sum_{i=1}^{M}f_{i}(x_{k})*\hat{\bm{v}}_{i}\right),$ (2)
where $\hat{\bm{v}}_{i}$ is the lead eigenvector of the covariance matrix obtained from the predictions of a set of binary classifiers on a given query; in our case, these correspond to each user’s binary preference on a given pairwise comparison query. As shown in Parisi et al. (2014), crowd labels derived via the SML are typically more accurate than labels from any individual crowd member or from a majority vote, as unlike other methods such as majority vote, the SML approximates a maximum likelihood estimate of the labels. In addition, Parisi et al. (2014) showed that $\hat{\bm{v}}_{i}$ can rank the users by their estimated error rates, which approximate the ground truth error rates. We denote labels estimated by SML as $\hat{\bm{y}}_{SML}$.
Parisi et al. (2014) also analyze the performance of SML in the presence of a cartel, a small group of users who attempt to steer the overall ensemble solution away from the (unknown) ground truth, and proved that (1) SML is robust to the presence of cartels, and that (2) the weights of the eigenvector $\hat{\bm{v}}_{i}$ can be used to identify the presence of cartels in the crowd. Here, we re-define the notion of a cartel as a minority group who provides feedback that consistently differs from the majority, for example users who provide feedback towards goals that are different than the goals of the majority.
3.3 Crowdsourced Preference Learning for RL
Crowd-PrefRL leverages the crowd-aggregated preference labels $\hat{\bm{y}}=\hat{\bm{y}}_{SML}$ for all instances in $\mathcal{S}$. We denote via $\mathcal{D}$ the dataset of all pairs $\{(x,y)\}$, where $x\in\mathcal{S}$ and $y$ is the corresponding crowd-aggregated label in $\hat{\bm{y}}$. To learn useful behaviors given these aggregated labels, we assume that the majority of users (though not necessarily all users) have a shared understanding of desired behavior, which can be quantified in terms of an underlying reward function. Given the crowd-aggregated labels, our objective is to learn the reward function, and ultimately a policy, that preserves the underlying user preferences Christiano et al. (2017). This is done in two steps: (1) a reward function learning step, where we optimize the learned reward via supervised learning from the preference labels $\hat{\bm{y}}$ over segment pairs generated by the policy $\pi$, and (2) a policy learning step, where the agent performs rollouts in the environment and optimizes its policy against the current reward function estimate. Segment pairs for preference queries are selected according to an entropy-based measure from an ensemble of reward predictors, similarly to Lee et al. (2021). These steps are then repeated for a set number of environment steps. Algorithm 1 shows the algorithmic steps needed for preference-based RL with learned rewards from crowds, and is analogous to Lee et al. (2021) with red text denoting the modifications needed to crowdsource the preference labels rather than obtaining them from a single human.
Similarly to Lee et al. (2021); Christiano et al. (2017), we learn a reward model $\widehat{r}_{\psi}$, modeled as a neural network with parameters $\psi$, that predicts the user preferences as follows:
$P_{\psi}[A\succ\ B]=\frac{\exp\sum_{t}\widehat{r}_{\psi}(\mathbf{s}_{t}^{A},\mathbf{a}_{t}^{A})}{\sum_{i\in\{A,B\}}\exp\sum_{t}\widehat{r}_{\psi}(\mathbf{s}_{t}^{i},\mathbf{a}_{t}^{i})}.$ (3)
The reward function $\widehat{r}_{\psi}$ is then updated by minimizing the following cross-entropy loss:
$\mathcal{L}^{\tt Reward}=-\operatorname*{\mathbb{E}}_{(A,B,y)\sim\mathcal{D}}\Big{[}$ $\mathbbm{1}_{[y=-1]}\log P_{\psi}[B\succ A]+$
$\mathbbm{1}_{[y=1]}\log P_{\psi}[A\succ B]\Big{]},$ (4)
where $\mathbbm{1}_{[\cdot]}$ denotes an indicator variable. Given the learned reward function $\widehat{r}_{\psi}$, a policy $\pi$ can be learned using any RL algorithm. In this work, we use PrefPPO Lee et al. (2021), instantiating Crowd-PrefRL as Crowd-PrefPPO. PrefPPO uses Proximal Policy Optimization (PPO), a state-of-the-art on-policy RL algorithm Schulman et al. (2017). As recommended in Lee et al. (2021), we use on-policy methods to reduce the effects of non-stationarity induced by learning a reward function and policy simultaneously.
4 Experiments
4.1 Experiment Setup
We conduct experiments to evaluate the proposed approach to crowdsourcing human feedback for preference learning. Our experiments conduct RL from human preferences using the PrefPPO implementation in Lee et al. (2021). These experiments use the Walker-walk, Quadruped-walk, and Cheetah-run environments from the DMControl suite Tassa et al. (2018), as two of these environments (Walker-walk and Quadruped-walk) were previously used in Lee et al. (2021). For each environment, we report results for Crowd-PrefPPO (SML preference estimation) and the following comparisons: Crowd-PrefPPO-MAJ (in which the majority vote is used instead of the SML estimate) and Crowd-PrefPPO-Oracle (in which the preference is given to the segment with the highest ground-truth environment reward); in addition, we compare Crowd-PrefPPO’s estimated preference labels to those from the best member of each crowd. We perform 10 total runs of each algorithm comparison in each environment for 4M environment steps. Methods are compared via the mean and standard error over 6 runs, where the top 2 and bottom 2 runs are omitted (measured by episode return at 4M environment steps) to reduce the impact of outliers (similar to the InterQuartile Mean, IQM Agarwal et al. (2021)). All remaining hyper-parameters for our method can be found in the Appendix, and notably, are the same as reported in Lee et al. (2021).
4.2 Results
Figure 2: Difference in MAJ and SML preference prediction error rates for different levels of variability (standard deviation) in user error rates. Values are calculated across 100 randomly-sampled crowds of 7, 11 and 15 simulated users for the Walker-walk environment. The horizontal dashed line at $y=0$ indicates where the MAJ and SML error rates are the same; points above $y=0$ indicate that the SML labels have lower error than the MAJ labels. We see a strong positive correlation, with increasing standard deviations (indicating a more diverse crowd) leading to improved SML label accuracy (lower error rate) compared to MAJ labels. In addition, we see a similar trend across all crowd sizes, suggesting that our method is robust across different sizes of crowds. Red dots indicate where SML outperforms the best crowd member, while blue dots indicate where the best crowd member outperforms SML.
We first study the effect of crowd size on the performance of crowdsourced preference learning by simulating 100 different random crowd configurations at each of three different crowd sizes ($M\in\{7,11,15\}$) according to the Stochastic Preference Model in Equation (1) for the Walker-Walk environment; see Methods for how these diverse crowds are simulated. Figure 2 depicts the difference in errors between the MAJ and SML preference estimates versus the diversity in users’ error rates with respect to the ground-truth reward (quantified via the standard deviation of user error across the crowd) for each simulated crowd size. To calculate these errors for each simulated crowd configuration, we first perform one full training run of Crowd-PrefPPO using Oracle preference labels to generate the full set of segment pairs needed for reward labeling. Then, we use these segment pairs and their ground-truth labels to construct the error rates for all users in the crowd and the error rates corresponding to the MAJ and SML labels. We note that results were similar for the other environments tested, and therefore show results for only the Walker-walk environment due to space constraints. There are several results here that are worth discussing. First, we see a strong positive correlation between the error rates of users in the crowd and the performance of SML compared to MAJ. In particular, as the user error increases (indicating a more diverse crowd), the better the SML labels perform compared to the MAJ labels. This indicates that the SML is effectively filtering out inconsistent preference feedback from users who have higher error rates and aggregating crowd decisions more consistently compared to MAJ. Second, we see that SML outperforms MAJ in nearly all cases ($\approx 90\%$ of points above $y=0$), suggesting that SML captures the crowd preferences more accurately than MAJ. Third, we see that the SML labels have lower errors than the best crowd member’s labels in nearly all cases (red dots). As the crowd size increases, we see fewer instances of SML performing worse than the best member of the crowd (blue dots), while simultaneously having more instances where SML outperforms MAJ (dots above $y=0$ dashed line). Finally, across all crowd sizes, the spread of the difference in MAJ errors and SML errors is fairly consistent (most points fall between $[0,6]\%$), suggesting that SML labels can outperform MAJ labels with most crowd configurations. While this improvement may seem small, it occurs during every reward learning iteration; we believe that this leads to compounding benefits during reward learning, and consequently, during agent policy learning.
Figure 3: (Top row) Comparison of Crowd-PrefPPO training curves with SML, MAJ and Oracle labeling across two different crowd configurations: [(a), left] a crowd for which SML is expected to outperform MAJ, and [(b), right] a crowd for which SML is expected to perform similarly to MAJ across the three different environments (Walker-walk, Quadruped-walk and Cheetah-run). (Bottom row) Comparison of MAJ and SML label prediction errors at each feedback iteration. Each plot shows the mean $\pm$ standard error of 6 out of 10 runs (the top and bottom 2 runs are omitted to reduce the effect of outliers, as detailed in the Experiment Setup).
Having studied how crowd configuration variability affects the respective prediction errors of the SML and MAJ crowd preference estimators, we next conduct experiments with Crowd-PrefPPO using the crowdsourced preference labels; in these experiments, we aim to determine if the improved label error from SML observed in Figure 2 impacts agent performance in a meaningful way. For this analysis, we select a fixed crowd size of $M=11$ users and sample two different crowd configurations: (1) Figure 3a studies a crowd for which SML outperforms MAJ by approx. $3\%$ in prediction error (following the analysis in Figure 2), and (2) Figure 3b studies a crowd for which SML performs similarly to MAJ in prediction error (similarly from the analysis in Figure 2). In addition, in the bottom row of Figure 3, we plot the error rates of the MAJ and SML label predictions relative to the ground-truth labels during each crowd feedback iteration for all three tested environments (200 queries per iteration for Quadruped-walk and 100 otherwise; see the Appendix for more details).
Figure 4: Plot of ground-truth error rate for each user in the crowd vs. the SML weight vector $\bm{v}_{i}$ for the same two SML configurations shown in Figure 3 in the Walker-walk environment. Across both conditions, we see a significant correlation between the weights generated by the SML and individual crowd member performance measured by the users’ (unknown) error rates, suggesting that users can be ranked accurately in the absence of ground-truth error information about the crowd.
In Figure 3a, we see that Crowd-PrefPPO outperforms Crowd-PrefPPO-MAJ in terms of trajectory return for all tested environments, with Crowd-PrefPPO achieving close to Oracle performance for Walker-walk and Quadruped-walk. In addition, the SML label errors are lower than the MAJ label errors at each feedback iteration in most cases for each environment. The lower SML label errors imply that there is a compounding benefit to using SML labels rather than MAJ labels for reward learning; in early feedback iterations, the learned reward function has lower error and thus is closer to the Oracle reward function compared to MAJ. This improves overall agent training, as (1) the agent receives more accurate rewards for its actions, enabling it to learn better behaviors earlier in the training loop, and (2) future crowd feedback iterations receive “better” segment pairs for the crowd to label, yielding more useful information about the reward function. We speculate that this is why Crowd-PrefPPO’s performance compares fairly well with that of Crowd-PrefPPO-Oracle. We see a similar, but weaker, trend in Figure 3b, where SML is expected to yield a similar label accuracy to MAJ for the given crowd configuration. There is generally less spread in episode returns between SML and MAJ for the three tested environments (top row of Figure 3b), and we see that the crowd feedback errors are closer between the methods during all crowd feedback iterations (bottom row of Figure 3b), indicating only minimal benefits of SML labels over MAJ labels. However, we do see that Crowd-PrefPPO can sometimes perform quite well overall, for example in Quadruped-walk.
Furthermore, we analyze the crowd member ranking derived by SML for the same two conditions shown in Figure 3 and the Walker-walk environment (results were similar for all environments, thus we report only one due to space constraints). While Crowd-PrefPPO calculates the eigenvector $\bm{v}_{i}$ during each feedback iteration, we here instead calculate $\bm{v}_{i}$ using all crowdsourced preference labels for simplicity of visualization. Results are shown in Figure 4. We see that across both conditions, the SML-derived member rankings are significantly correlated to the individual members’ error rates (i.e., higher SML weight values indicate lower error rates), suggesting that even when SML performs similarly to MAJ in label error (Figure 4b), SML can still identify the best members in the crowd. This showcases additional utility of SML aggregation over MAJ aggregation even in the situation where SML provides no benefit in label accuracy compared to MAJ.
Figure 5: Plot of ground-truth error rate vs. the SML weight vector $\bm{v}_{i}$ for a crowd configuration where there are three distinct groups of users: users who provide feedback according to (1) the ground-truth reward $r$ (blue), (2) at random (black), and (3) the opposite of the ground-truth reward, $-r$ (red). We see distinct groupings of users by their feedback category and a strong correlation between their (unknown) error rates and the SML weight vector $\bm{v}_{i}$, suggesting that minority groups of users can be accurately identified.
Finally, we test the ability of SML to detect the presence of minority group feedback based purely on the preference labels. We follow the same analysis used in Figure 4, except that we simulate a crowd of 15 users according to three groups: users who provide noisy preference feedback according to (1) the ground-truth reward $r$ (9 users), (2) at random (3 users), and (3) $-r$, the opposite of the ground-truth (3 users). Results are shown in Figure 5. As expected, we see that MAJ error rates are significantly higher compared to SML rates, since MAJ is more sensitive to the presence of noisy labels. In addition, we see distinct clusters of users based on their feedback category, suggesting that the SML weights $\bm{v}_{i}$ can be used to identify the presence of minority groups of users, as well as the identities of the minority group members. These results suggest that Crowd-PrefRL could be used to cluster users via their objectives and to learn reward functions and corresponding behaviors specific to each user cluster. Such functionality could enable a behavior-based characterization of the objective associated with each user cluster.
5 Discussion
This study aimed to determine how to (1) effectively combine preference feedback from crowds of unknown expertise and reliability in the absence of ground-truth information, (2) identify the “best” users in a given crowd in an unsupervised manner, (3) learn a crowdsourced reward function that is more accurate than the reward learned from any individual user, and (4) identify minority groups based purely on their preference feedback. We propose the Crowd-PrefRL framework and instantiate it as Crowd-PrefPPO. Our evaluation of Crowd-PrefPPO indicates that we can use methods derived from unsupervised ensemble learning (SML, Parisi et al. (2014)) to effectively aggregate user preference feedback across diverse crowds in the absence of any ground-truth information about the users. Furthermore, our evaluation suggests that Crowd-PrefPPO can identify which users in a crowd are likely most reliable and that the reward function learned from crowd-aggregated labels is more accurate than the reward derived from any single user’s feedback (even the best user!). Finally, we show the possibility of detecting clusters of minority feedback in an unsupervised manner, potentially enabling a deeper understanding of model behavior (e.g. detecting the presence of multiple user viewpoints, characterizing the associated objectives, and mitigating effects of malicious user data Siththaranjan, Laidlaw, and Hadfield-Menell (2023); Ouyang et al. (2022)). To the best of our knowledge, this is the first demonstration of (1) learning crowd-level preference labels for preference-based RL via unsupervised ensemble learning methods in the online RL setting and (2) identification of minority groups in the crowd based purely on the crowd’s preference feedback.
Previous work has studied crowdsourcing preference labels for reward learning in the offline RL case  Zhang and Kashima (2023), as opposed to the online RL case we study here. In the offline case, a dataset of preference pairs is already provided a priori, and the goal is to learn a reward function from this static dataset. This limits the expressiveness of the learned reward function to the behaviors contained in the dataset, whereas in our work, the preference pairs for crowd labeling are sampled throughout agent training, and by extension the learned reward function continuously improves with better agent performance and better environment exploration. In addition, the method in  Zhang and Kashima (2023) requires training a separate model (e.g. fully connected network) to infer crowd labels and member reliabilities, requiring additional model training and tuning and increased data requirements (50K preference pairs from 2,500 crowdworkers). In contrast, our approach to inferring crowd labels by the SML is learning and optimization-free: we need no hyperparameters to specify, no model to train, and no optimization procedure to learn a crowd label. Thus, we believe it is more suitable in the online RL case where the number of labeled preference pairs is small compared to the offline RL case due to practical constraints (agent is training during human feedback collection). Our work enables crowds to update the reward function throughout agent training using preference feedback on just a handful of trials and furthermore, allows for the detection of minority feedback groups, which has important modeling applications in various domains.
We showed that SML labels outperform MAJ labels in nearly all cases across different crowd sizes (Figure 2) and different tasks (Figure 3). Calculating SML weights is computationally very fast, only requiring a few lines of code in modern programming languages (e.g. Python). Given the increasing popularity of crowdsourcing platforms for data collection in human-interactive RL studies (e.g. Amazon Mechanical Turk), these results indicate that simple code modifications can result in significant performance gains. For example, Brown et al. (2019) used a crowdsourcing platform for their human subjects data analysis and estimated preference labels for segments by simply taking a majority vote over multiple human users. It would be interesting to apply Crowd-PrefRL to aggregate crowd preferences in such existing studies to determine the extent to which Crowd-PrefRL improves over results from a majority vote or offline RL.
We believe that Crowd-PrefRL could be combined with recent advances that make preference-based RL methods more sample-efficient Lee, Smith, and Abbeel (2021); Park et al. (2022). Improvements to sample efficiency are made primarily at the algorithmic or data level, as opposed to modifying the labels used to learn the reward. For example, PEBBLE Lee, Smith, and Abbeel (2021) combines unsupervised pre-training with off-policy re-labeling of past data to achieve improved sample efficiency compared to PrefPPO. SURF Park et al. (2022) uses semi-supervised learning methods that combine data augmentation approaches with confidence-based pseudo-labels of unlabeled experiences to train the preference predictor. Combining Crowd-PrefRL with these orthogonal advancements could potentially lead to further improvements in RL agent training performance.
Several downsides to our approach are important to discuss. First, each user in the crowd must provide a preference label for each preference query. This results in significant human labeling cost if implemented in real-world studies. Our results indicate that Crowd-PrefPPO can work well in fairly small crowds with relatively small numbers of labels (100 queries per user per reward learning iteration). Future work will focus on the possibility of using Crowd-PrefRL when user feedback is sparse (i.e., not every user provides a preference for every segment pair). In addition, this study uses only simulated human feedback and does not include human-subjects data collection. Rather, we simulate human feedback via the Stochastic Preference Model, which allows for a thorough and systematic evaluation of Crowd-PrefRL and an analysis of its properties in a highly controlled setting. Future work will evaluate this method in a human subjects study.
Appendix
Experiment Setup Hyper-parameters
Our experiments use the Walker-walk, Quadruped-walk, and Cheetah-run environments from the DMControl suite Tassa et al. (2018), as two of these environments (Walker-walk and Quadruped-walk) were previously used in Lee et al. (2021). All hyper-parameters for our method can be found in Table 1, and notably, are the same as reported in Lee et al. (2021).
Hyperparameter Value
GAE parameter $\lambda$ $0.9$ (Quadruped), $0.92$ (otherwise)
Hidden units per each layer $256$
Segment of length $50$
# of layers $3$
Learning rate $0.00005$
Batch Size $128$ (Quadruped), $64$ (otherwise)
Discount ${\bar{\gamma}}$99.99.99.99
Frequency of feedback $32000$
# of envs per worker $16$ (Quadruped), $32$ (otherwise)
PPO clip range $0.4$
Entropy bonus $0.0$
# of timesteps per rollout $500$
Maximum budget $2000$ (Quadruped), $1000$ (otherwise)
# of feedbacks per session $200$ (Quadruped), $100$ (otherwise)
Table 1: Hyper-parameters of the Crowd-PrefPPO algorithm for all environments.
Ethical Statement
In any real-world crowdsourcing application, it will be important to protect user privacy such that only anonymized user data is stored and so that user data is only stored given user consent. Furthermore, systems trained via crowdsourced data must be able to recognize and account for differing user perspectives. Critically, in applications from household robotics to autonomous driving to LLM-based assistants, users must not be forced into options optimized for other users; rather, systems should utilize crowd-aggregated information as a baseline from which to accelerate and improve downstream personalization to specific individual users. Meanwhile, it is crucial to identify and reject malicious user data to avoid biasing models toward toxic or other undesirable behaviors. We believe that Crowd-PrefRL can help to facilitate such system capabilities by identifying minority viewpoints, not only learning a policy consistent with the majority viewpoint, but also enabling developers to model behaviors corresponding to distinct user clusters to improve overall modeling. Finally, when analyzing different user objectives, it will be important to consider the potential implications of empowering a few individuals to decide which user objectives are worthwhile, toxic, etc.
Acknowledgments
This project was sponsored by the DEVCOM Army Research Laboratory. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.